Sarah and Javier are the partners in a local coffee shop. They needed $21,000 to start the business. They invested in the ratio 2:3, Sarah to Javier. How much did each invest?
Answer by ikleyn(52852)

.

    2x + 3x = 21000


    5x = 21000

     x = 21000/5 = 4200.


ANSWER.  Sarah invested  2*4200 = 8400 dollars.  Javier invested  3*4200 = 12600 dollars.

Solved.